You are going to deposit $3,000 in an account that pays .36 percent interest compounded monthly. How much will you have in 6 years?
Answer #1

PV =3000
Rate per month =0.36%
Number of Periods =12*6 =72
FV =PV*(1+r)^n =3000*(1+0.36%)^72=3885.87